Simon Law

8–4137 Av. de l’Esplanade., Montréal, QC. H2W 1S9. (514) 839-4887 sfllaw@sfllaw.ca

Soft skills:

• Speaks English natively. Adores the written word.

• Able to form structured and convincing logical arguments.

• Facilitates groups well, encouraging creative ideas and solid decisions.

• Excellent presentation skills. Fluent technical writer.

• Loves to teach, both in a classroom and tutor-student setting.

Application programming:

• Application languages: Python, C++ with STL, C

• Scripting languages: JavaScript, Bourne shell, sed & awk, Perl

• Macro languages: M4, TEX, L^{A}TEX

• Markup languages: XHTML with CSS, XML

• Declarative languages: SQL 99, GNU make

• Functional languages: Emacs Lisp

System administration:

• Proficient with Ubuntu and Debian GNU/Linux system administration.

• Proficient with Amazon Web Services (EC2, S3)

• Proficient with Mercurial, Buildbot, FogBugz, TWiki

• Familiar with Mac OS X.

Network administration:

• Familiar with Intrusion Detection Systems and Vulnerability Assessments. (nMap, Snort, Nessus)

• Implemented Virtual Private Networking solutions. (PPTP, L2TP, SSH)

• Understands TCP/IP suite of protocols. Includes proficiency with routing, packet filtering, and proxying.

Work Experience:

Akoha Inc.; Drunken Master; Montréal April 2007 – Present • Co-designing the Akoha website, from database schemas to user experience.

• Developing the Akoha website, from the Django backend to the web frontend.

• Building an automated staging and deployment system for the Akoha website.

• Contributes to the Django web framework, both to the core software and third-party modules.

• Maintains the Mercurial Forest extension, which allows nested repositories to act as one.

• Built the core development infrastructure, including bug tracking, wikis, source control, build system, automated testing, and continuous integration.

• Managed customer service and fulfillment issues.

Canonical Ltd.; Quality Controller; Montréal April 2006 – February 2007 • Started the Quality Assurance department for Ubuntu 7.04 (Feisty Fawn)

• Developed training materials and policies for triaging bug reports for a large GNU/Linux distribution.

• Ran online seminars and managed the IRC channels to build a team of enthusiastic volunteers who triaged over 300 bug reports per day.

• Co-ordinated with Customer Support and Hardware Certification teams to escalate key issues to Development.

• Co-designed apport, a generic, automated crash reporting tool that collects debugging information from users’ systems.

Net Integration Technologies; Drunken Master; Montréal September 2003 – April 2006 • Acted as Release Manager for our Free Software offerings. Packages our software for the Debian Project.

• Acted as Project Manager for Nitix for SAGE Accpac ERP.

• Advocated products and technologies at various conferences and organisations.

• Coded network and user-interface software in C++ using the WvStreams library for the Net Integrator product.

• Developed and maintained UniConf, a universal configuration system.

• Architected and designed a clustered, automated test harness for the Nitix operating system.

Maplesoft, a division of Waterloo Maple Inc.; Toolsmith; Waterloo January 2003 – April 2003 • Maintained scripts to build Maple on 8 different platforms.

• Designed and implemented an end-to-end build system in Apache Ant.

Bank of Montreal; IS Security Analyst; Toronto September 2001 – August 2002 • Helped design a network security appliance merging Debian, nMap, Snort, ACID, Nessus, PostgreSQL, SSH, Netfilter.

• Wireless (802.11b) vulnerability assessments and external penetration testing.

• Designed a data warehouse for reporting purposes, and did a first phase implementation.

Education:

University of Waterloo, Waterloo September 1999 – April 2004 • B.ASc. in Computer Engineering with an Option in Software Engineering.

EWA–Canada CanCERT, Ottawa September 2001 • Completed an Intrusion Detection Systems and Incident Response course.

Random facts:

• Worst vice: bad puns

• Best vice: throwing dinner parties

• Hobbies: teaching, cooking, eating, photography, attending concerts, coding, reading, electronics, writing, mail and news.

• I am a Free Software advocate and a GNU contributor.

• I have a meagre set of web pages at http://sfllaw.ca/~sfllaw/

• I have a not-so-meagre weblog at http://sfllaw.livejournal.com/
• I have a volumnous photo album at http://flickr.com/photos/sfllaw/
• Stalk me online at: http://google.com/search?q=sfllaw